X company currently has outstanding 1,00,000 shares selling at rs 100 each .the firm is thinking of declaring a dividend rs 5 per shareat the end of the current year. The capitalization of this type of firm is 10%. What will be the price of the share at the end of the year if dividend is declared.
1
Expert's answer
2021-09-08T09:36:56-0400

"P_0 =\frac{(D_1 + P_1)} {1+ K_0}"

"P_1 = P_0 (1+K_0)-D_1"



"P_0=" Current market price of the share "=rs 1000"

"D_1=" Expected dividend at the end of year one "=rs 5"

"K_0=" Cost of capital/equity"= 10\%"

"P_1=" Expected price of the share at the end of the year"=?"


"P_1= 100(1+0.10)-5"

"=110-5"

"=rs 105"


Therefore, if the dividend is declared the price of the share at the end of the year will be rs 105